# Benchmarking LLMs’ Judgments with No Gold Standard

### v1

@@ 4.3. Result 3: Better Robustness against Manipulation
Manipulation Strategies
We now introduce the manipulation strategies. The detailed setup, including the prompts and examples, is provided in Appendix D.2.
  1. (1)
@@
Figure 4. Meaningless Elongation
Statistics and Results
The SMDs shown in Tabel 2 demonstrate that LMexaminer has significant score increases after all three manipulations, indicating its lack of manipulation-resistance. In contrast, GEM and GEM-S exhibit no significant score increase across these manipulations. Furthermore, comparing SMDs after degradations and manipulations in Table 2, we observe that the scores of GEM and GEM-S decrease less after manipulations compared to degradations, indicating that manipulations result in less information loss than degradations.
5. GRE-bench: Benchmarking LLMs with GEM
@@
Figure 5. Results of GRE-bench based on three evaluation metrics with 90% confidence intervals vs. model parameter size. The grey line represents the average human baseline, with the 90% confidence interval shaded in grey.
@@
Notably, several models with large parameter sizes surpass the human baseline (gray line) when using GEM and GEM-S (abstract) metrics. We hypothesize that the reason is that strong LLMs like Claude-3 opus can effectively retrieve contributions and limitations stated by the authors in the paper, and will include that in their reviews. In contrast, human reviewers tend to omit some judgments that are less important or already stated by the author due to writing and expression costs, resulting in a loss of mutual information. This hypothesis can be supported by the observation that, when we condition out the author-stated strengths and weaknesses, the human reviewer has the highest GRE-bench score based on GEM-S (ASSW).
This observation further validates the hierarchical information structure as we discussed in Section 3.2. By conditioning out different levels of information, the different benchmarks capture subtly different abilities. GEM-S (ASSW) focuses more on quality of critical judgment, while GEM and GEM-S (abstract) are more effective in evaluating LLMs’ abilities in more basic aspects, including retrieving useful information from the paper, which is especially valuable for evaluating LLMs with smaller parameter size.
@@
We present the Spearman’s correlation coefficients between the GRE-bench scores of all the models and their corresponding scores on other benchmarks in Table 3, providing insights into which factors most strongly influence review generation quality. Notably, GRE-bench highly correlates with HellaSWAG and ARC-C which measures reasoning ability, and less correlates with benchmarks for coding ability (HumanEval) or math ability (MATH, GSM8K).
@@
Table 3. Spearman’s correlation coefficients between GRE-bench and other popular benchmarks.
Consistency over Various Model Sizes
The Spearman’s correlation coefficients between the GRE-bench scores based on GEM-S (ASSW) metric computed using Llama-3.1 8B and that same score computed with Llama-3.1 70B is 0.93. This indicates that the score is largely robust with respect to the specific evaluation LM used. The analogous coefficient for the GRE-bench scores based on GEM-S (abstract) and GEM are 0.92 and 0.90 respectively. We further provide the comprehensive results of GRE-bench with Llama-3.1 8B in Table 7 of Appendix A.3. This consistency shows that with our metrics, a small model with 8B parameters that can be deployed on a single NVIDIA L4 GPU, plus proper preprocessing, is still effective in evaluating the quality of judgments generated by much larger models, though larger evaluation LM may have better evaluation performance.
6. Conclusion and Discussion
@@
We list several potential directions for future work. We can use new open-access peer review data to update GRE-bench every year and observe how the evaluation evolves over time, to monitor any improvements in LLMs’ peer review capabilities. This will also ensure the benchmark stays current and minimizes data contamination.
Moreover, for specific tasks with a rich dataset, e.g., peer review in ICLR, we can finetune the evaluation-LM to better estimate the joint distribution. However, the widely-used supervised finetuning may not be suitable for subjective tasks, as the responses in the dataset may have different judgments on the same task. We provide initial results in Appendix E, which show better efficacy than baselines, but are not significantly better than the un-finetuned model, thus, a more suitable finetuning method may be further explored.
@@ Acknowledgment
We would like to thank Sanzeed Anwar from University of Michigan for data processing. We would also like to thank Zebang Xu and Tianzi Liu from Cornell University for consultation on the statistical methods applied in our experiments.
@@ Appendix A Additional Results
A.1. Result 2 and 3 with Confidence Intervals
@@
Table 4. Standardized Mean Differences (SMD) $d:=(\mu^{\prime}-\mu)/{\sigma_{\text{pooled}}}$ of scores after degradations of various evaluation metrics with 95% confidence intervals in parentheses. Significant score decreases ($p<0.05$) after degradations are highlighted in bold green, implying the evaluation metric can effectively penalize the degradation.
@@
Table 5. Standardized Mean Differences (SMD) of scores $d:=(\mu^{\prime}-\mu)/{\sigma_{\text{pooled}}}$ after manipulations of various evaluation metrics with 95% confidence intervals in parentheses. Significant score increase ($p<0.05$) after manipulations are highlighted in bold red, implying the evaluation metric are not robust against the manipulation.
A.2. Detailed Results of GRE-bench with Llama3.1 70B
We present numerical results of various language models on the GRE-bench based on GEM, GEM-S (abstract), and GEM-S (ASSW) with Llama3.1 70B, alongside their estimated parameter sizes in Table 6. The corresponding results of GRE-bench with Llama3.1 8B is shown in Table 7.
@@
Table 6. Results of GRE-bench based on GEM, GEM-S (abstract), and GEM-S (ASSW) with Llama3.1 70B. In the column of parameter size, the symbol $\sim$ indicates that this size is estimated.
A.3. Detailed Results of GRE-bench with Llama3.1 8B
@@
Table 7. Results of GRE-bench based on GEM, GEM-S (abstract), and GEM-S (ASSW) with Llama3.1 8B. In the column of parameter size, the symbol $\sim$ indicates that this size is estimated.

### v2

@@ 4.3 Result 3: Better Robustness against Manipulation
Manipulation Strategies. We now introduce the manipulation strategies. The detailed setup, including the prompts and examples, is provided in Appendix D.4.
  1. 1.
@@
Figure 4: Meaningless Elongation
Statistics and Results. The SMDs shown in Table 2 demonstrate that LMexaminer has significant score increases after all three manipulations, indicating its lack of manipulation-resistance. In contrast, GEM and GEM-S exhibit no significant score increase across these manipulations. Furthermore, comparing SMDs after degradations and manipulations in Table 2, we observe that the scores of GEM and GEM-S decrease less after manipulations compared to degradations, indicating that manipulations result in less information loss than degradations.
5 GRE-bench: Benchmarking LLMs with GEM
@@
Figure 5: Results of GRE-bench based on three evaluation metrics with 90% confidence intervals vs. model parameter sizes [^6] In Figure 6, the parameter sizes of closed-source models are estimated from unofficial sources. This information is only to provide context, and none of our conclusions or key findings depend on these estimates.indicated by color. The grey line represents the average human baseline, with the 90% confidence interval shaded in grey.
@@
Notably, several models with large parameter sizes surpass the human baseline (gray line) when using GEM and GEM-S (abstract) metrics. We hypothesize this is because strong LLMs like Claude-3 opus can effectively retrieve contributions and limitations stated by the authors, and will include that in their reviews. In contrast, human reviewers tend to omit judgments that are less important or already stated by the authors due to writing costs, resulting in a loss of mutual information. This hypothesis is supported by the observation that, when we condition out the author-stated strengths and weaknesses, the human reviewer has the highest GRE-bench score based on GEM-S (ASSW).
This observation further validates the hierarchical information structure discussed in Section 3.2. By conditioning out different levels of information, the benchmark captures subtly different abilities. GEM-S (ASSW) emphasizes the quality of critical judgment, while GEM and GEM-S (abstract) focus more on evaluating LLMs’ foundational abilities, including retrieving useful information from the paper, which is especially valuable for differentiating LLMs with smaller parameter size.
@@
We present the Spearman’s correlation coefficients between the GRE-bench scores of all the models and their corresponding scores on other benchmarks [^7] We collect benchmark data from online sources, such as technical reports of language models. Since some models have missing results for specific benchmarks, we exclude those entries when computing correlation. For more details, please refer to our code repository. in Table 6 in Appendix A.2, providing insights into which factors most strongly influence review generation quality. Notably, GRE-bench highly correlates with HellaSWAG and ARC-C which measures reasoning ability, and less correlates with benchmarks for coding ability (HumanEval) or math ability (MATH, GSM8K).
@@
Table 3: Spearman’s correlation coefficients between GRE-bench and other popular benchmarks.
Consistent results over various Evaluation LMs.
The Spearman’s correlation coefficients between the GRE-bench scores based on GEM metric computed using Llama-3.1 8B and that same score computed with Llama-3.1 70B is 0.93. This indicates that the score is largely robust with respect to the specific evaluation LM used. The analogous coefficient for the GRE-bench scores based on GEM-S (abstract) and GEM-S (ASSW) are 0.92 and 0.90 respectively. We further provide the comprehensive results of GRE-bench with Llama-3.1 8B in Table 8 in Appendix A.2. This consistency shows that with our metrics, a small model with 8B parameters that can be deployed on a single NVIDIA L4 GPU, plus proper preprocessing, is still effective in evaluating the quality of judgments generated by much larger models, though larger evaluation LM may have better evaluation performance.
6 Conclusion and Discussion
@@
For future work, we can use new open-access peer review data to update GRE-bench every year and observe how the evaluation evolves over time, to monitor any improvements in LLMs’ peer review capabilities. This will also ensure the benchmark stays current and minimizes data contamination. Additionally, exploring GEM’s potential to evaluate subjective content in other scenarios, such as book reviews and movie reviews, could broaden its applicability and further demonstrate its versatility. We provide initial results in validating GEM on a Yelp restaurant review dataset in Appendix A.4.
@@ Acknowledgment
We would like to thank Sanzeed Anwar from University of Michigan for data processing. We also thank Zebang Xu and Tianzi Liu from Cornell University for their valuable consultation on the statistical methods employed in our experiments. Additionally, we appreciate the insightful feedback and suggestions provided by the anonymous reviewers at ICLR 2025.
Reproducibility Statement
In this section, we demonstrate the efforts we have made to ensure better reproducibility of the paper.
Complete Theoretical Proof
In Appendix B, we provide a full proof of the conclusions claimed in the paper.
Ready-to-Use Code
We have made all the code for data collection, model fine-tuning, metric computation, and experiments available on GitHub. The repository is at https://github.com/yx-lu/Benchmarking-LLMs--Judgments-with-No-Gold-Standard. Additionally, to address potential hardware issues for reproducing the results, our code can be deployed on Google Colab (requiring a T4 or higher GPU, with A100 being recommended). Experiments with Llama-3.1-8B as the evaluation LM are conducted on Google Colab instances with NVIDIA A100 GPU. Experiments in Section 5 employ Llama-3.1-70B as the evaluation LM requiring more GPU VRAM, thus, we use rental instances with a configuration of two NVIDIA H100 GPUs.
Clear Implementation and Experimental Details
In Appendix C, we thoroughly discuss the implementation of our GEM and GEM-S metrics. In Appendix D, we provide detailed explanations of the experimental procedures, including how we degrade and manipulate reviews, as well as the methodology for computing the baseline metrics.
@@ Appendix A Additional Results
In this section, we provide additional detailed results that were omitted from the main paper and present the results of additional experiments to further validate our GEM metrics.
A.1 Result 2 and 3 with Confidence Intervals
@@
Table 4: Standardized Mean Differences (SMD) $d:=(\mu^{\prime}-\mu)/{\sigma_{\text{pooled}}}$ of scores after degradations of various evaluation metrics with 95% confidence intervals in parentheses. Significant score decreases ($p<0.05$) after degradations are highlighted in bold green, implying the evaluation metric can effectively penalize the degradation.
@@
Table 5: Standardized Mean Differences (SMD) of scores $d:=(\mu^{\prime}-\mu)/{\sigma_{\text{pooled}}}$ after manipulations of various evaluation metrics with 95% confidence intervals in parentheses. Significant score increase ($p<0.05$) after manipulations are highlighted in bold red, implying the evaluation metric are not robust against the manipulation.
A.2 Detailed Results of GRE-bench
GRE-bench vs. Other benchmarks
We present the Spearman’s correlation coefficients between the GRE-bench scores of all the models and their corresponding scores on other benchmarks [^8] We collect benchmark data from online sources, such as technical reports of language models. Since some models have missing results for specific benchmarks, we exclude those entries when computing correlation. For more details, please refer to our code repository. in Table 6, providing insights into which factors most strongly influence review generation quality. Notably, GRE-bench highly correlates with HellaSWAG and ARC-C which measures reasoning ability, and less correlates with benchmarks for coding ability (HumanEval) or math ability (MATH, GSM8K).
Base Metric MMLU ARC-C HellaSwag GSM8K MATH HumanEval GPQA
GEM 0.55 0.68 0.74 0.58 0.37 0.36 0.60
GEM-S (abstract) 0.66 0.70 0.82 0.73 0.43 0.43 0.67
GEM-S (ASSW) 0.68 0.78 0.84 0.73 0.43 0.48 0.71
Table 6: Spearman’s correlation coefficients between GRE-bench and other popular benchmarks.
Consistent results over various Evaluation LMs.
The Spearman’s correlation coefficients between the GRE-bench scores based on GEM metric computed using Llama-3.1 8B and that same score computed with Llama-3.1 70B is 0.93. This indicates that the score is largely robust with respect to the specific evaluation LM used. The analogous coefficient for the GRE-bench scores based on GEM-S (abstract) and GEM-S (ASSW) are 0.92 and 0.90 respectively. We further provide the comprehensive results of GRE-bench with Llama-3.1 8B in Table 8. This consistency shows that with our metrics, a small model with 8B parameters that can be deployed on a single NVIDIA L4 GPU, plus proper preprocessing, is still effective in evaluating the quality of judgments generated by much larger models, though larger evaluation LM may have better evaluation performance.
Numerical results.
We present numerical results of various language models on the GRE-bench based on GEM, GEM-S (abstract), and GEM-S (ASSW) computed with Llama3.1 70B as the evaluation LM, alongside their estimated parameter sizes in Table 7. Note that some estimations of the parameter sizes are from unofficial sources. This information is only to provide context, and none of our conclusions depend on these estimates. The corresponding results of GRE-bench computed with Llama3.1 8B is shown in Table 8.
@@ A.2 Detailed Results of GRE-bench
Table 7: Results of GRE-bench based on GEM, GEM-S (abstract), and GEM-S (ASSW) with Llama3.1 70B. In the column of parameter size, the symbol $\sim$ indicates that this size is estimated. For mix-of-expert (MoE) models, we show the total parameter size instead of the activated parameter size.
@@
Table 8: Results of GRE-bench based on GEM, GEM-S (abstract), and GEM-S (ASSW) with Llama3.1 8B. In the column of parameter size, the symbol $\sim$ indicates that this size is estimated. For mix-of-expert (MoE) models, we show the total parameter size instead of activated parameter size.
A.3 Consistent Results over Preprocessing Models
To answer the question “Does using an open-sourced preprocessing model (rather than GPT4o) still lead to effective GEM metric and similar GRE-bench results?”, we use a Llama3.2-90B-Vision-Instruct model for preprocessing and re-run all validating experiments mentioned in Section 4. We find that most experimental results remain valid except for two manipulation tests. The GRE-bench results also remain highly consistent with the results of using GPT4o preprocessing.
Positive Correlation with Human Annotation
GEM still shows a significant positive correlation with human annotation in the proposal peer grading dataset (Section 4.1), however, the Spearman’s correlation coefficient between GEM scores and instructor grades drops from 0.433 (GPT4o preprocessing) to 0.274 (Llama3.2-90b preprocessing).
Sensitivity to Degradation
In the degradation experiment (Section 4.2), GEM with Llama3.2-90b preprocessing also shows significant sensitivity to all degradations and even performs better in “sentence deletion” than using GPT4o to preprocess.
GEM -0.566 -0.317 -0.078
(Llama Preproccessing) (-0.625,-0.507) (-0.371,-0.263) (-0.154,-0.002)
GEMS -0.549 -0.205 -0.327
(Llama Preproccessing) (-0.607,-0.491) (-0.258,-0.153) (-0.399,-0.254)
Table 9: Standardized Mean Differences (SMD) $d:=(\mu^{\prime}-\mu)/{\sigma_{\text{pooled}}}$ of scores after degradations of GEM and GEM-S metrics (preprocessing with Llama-3.2 90B) with 95% confidence intervals in parentheses. Significant score decreases ($p<0.05$) after degradations are highlighted in bold green, implying the evaluation metric can effectively penalize the degradation.
Weaker Robustness than GPT-4o Preprocessing
In the manipulation experiments (Section 4.2), GEM with Llama3.2-90b preprocessing is robust against “GPT4 rephrase” but fails in the “Llama-3.1 rephrase” and “Meaningless Elongation”.
GEM -0.132 0.124 0.138
(Llama Preproccessing) (-0.180,-0.084) (0.075,0.173) (0.086,0.189)
GEM-S -0.126 0.156 0.112
(Llama Preproccessing) (-0.173,-0.079) (0.109,0.203) (0.061,0.163)
Table 10: Standardized Mean Differences (SMD) $d:=(\mu^{\prime}-\mu)/{\sigma_{\text{pooled}}}$ of scores after manipulations of GEM and GEM-S metrics (preprocessing with Llama-3.2 90B) with 95% confidence intervals in parentheses. Significant score increases ($p<0.05$) after manipulations are highlighted in bold red, implying the evaluation metrics are not robust against the manipulation.
Consistent GRE-bench Results
Based on the GEM metric with llama3.2-90b preprocessing, we re-compute the GRE-bench scores with ICLR 2023 dataset (Section 5). The Spearman’s correlation coefficient between the GRE-bench scores based on the GEM metric with Llama3.2-90B preprocessing and the scores with GPT-4o preprocessing shown in Table 8 is 0.89. The analogous coefficient for the GRE-bench scores based on GEM-S (abstract) and GEM-S (ASSW) are 0.85 and 0.83 respectively. All detailed numerical results are provided in Table 11.
Model Name Para. Size GRE-bench
(in Billion) GEM GEM-S(abs.) GEM-S(ASSW)
openai/gpt-4o-mini (v 20240718) $\sim$ 8 45.76 19.58 18.27
openai/gpt-4o (v 20240513) $\sim$ 200 50.17 23.12 20.55
openai/gpt-3.5-turbo (v 20240125) $\sim$ 20 48.56 20.92 19.04
anthropic/claude-2 $\sim$ 137 54.61 26.07 23.93
anthropic/claude-3-haiku (v 20240307) $\sim$ 20 52.70 24.87 22.18
anthropic/claude-3-sonnet (v 20240229) $\sim$ 70 54.83 25.82 22.78
anthropic/claude-3-opus (v 20240229) $\sim$ 2000 62.08 33.13 30.53
google/gemini-pro (v 002) $\sim$ 20 46.75 21.33 20.43
google/gemini-pro-1.5 (v 002) $\sim$ 175 56.36 27.40 24.44
meta-llama/llama-3-8b-instruct 8 47.36 20.96 19.86
meta-llama/llama-3-70b-instruct 70 55.62 29.61 27.98
mistralai/mixtral-8x7b-instruct (v 0.1) 56 50.53 22.83 21.08
mistralai/mixtral-8x22b-instruct (v 0.1) 176 51.57 23.84 21.79
meta-llama/llama-3.1-8b-instruct 8 51.04 24.51 22.28
meta-llama/llama-3.1-70b-instruct 70 50.26 23.68 21.73
mistralai/mistral-large (v 2407) 123 51.81 23.59 21.20
mistralai/mistral-small (v 2409) 22 50.22 22.99 20.58
mistralai/mistral-tiny (v mistral-7b-0.3) 7 48.42 20.04 18.86
Table 11: Results of GRE-bench based on GEM, GEM-S (abstract), and GEM-S (ASSW) with Llama3.2 90B preprocessing. In the column of parameter size, the symbol $\sim$ indicates that this size is estimated.
Overall, as more advanced models can preprocess reports more accurately, the GEM metric will be more accurate and manipulation-resistant with superior preprocessing. Considering the frequent release and update of LLMs, utilizing more advanced LLMs for preprocessing enhances both the accuracy and robustness of the GEM metric. Consequently, our GEM metrics stand to benefit from the continuous advancements in LLM technology.
A.4 Validation on Yelp Restaurant Review Dataset
To further validate our GEM metrics’ effectiveness in evaluating judgments without gold-standard references in scenarios in addition to peer review. We use an open-access online business review data from Yelp, and repeat the same workflow in Section 4.2 and 4.3, to check how sensitive GEM metrics are to degradations and how robust GEM metrics are to manipulations.
The results are similar to those observed in experiments on ICLR dataset (Section 4.2 and 4.3). Our GEM and GEM-S demonstrates significant sensitivity to degradations and robustness against manipulations.
Yelp Review
We randomly sample 300 restaurants from the entire dataset which have at least 50 reviews with at least 1000 characters. For each restaurant, we randomly sample 3 customer reviews among all reviews with at least 1000 characters. Yelp reviews quality varies much more than ICLR reviews, thus, we use length as a heuristic filter to get reasonable quality reviews.
GEM Setup
We use GPT-4o for text preprocessing, and use Llama-3.1 8B as the evaluation-LM. For GEM-S, we use the categories of the restaurant as the synopsis, which is given by the Yelp dataset.
GEM demonstrates sensitivity to degradations.
We reuse the Sentence Deletion and Deletion & Completion degradation methods described in Section 4.2. However, since restaurant reviews lack “abstract”, the Abstract-only Review degradation cannot be applied. Instead, we propose Random Replacement degradation, where we degrade the original review $x_{i}$ by replacing it with a random review $x_{i}^{\prime}$ of another randomly selected restaurant.
Table 12 presents the standardized mean differences (SMD) with 95% confidence intervals for GEM and baseline metrics before and after these degradations. The results align with those observed in the ICLR dataset (Section 4.2): GEM and GEM-S consistently exhibit significant SMD across all three degradations. Additionally, while LMExaminer is more sensitive to degradations like Sentence Deletion and Deletion & Completion, which impair presentation quality, GEM shows greater sensitivity to the Random Replacement degradation, which severely compromises semantic informativeness.
Evaluation Metric Sentence Deletion Deletion & Completion Random Replacement
BLEU -0.348 -0.113 -0.231
(-0.436,-0.260) (-0.200,-0.026) (-0.355,-0.107)
ROUGE-L -0.677 0.009 -0.345
(-0.765,-0.589) (-0.064,0.081) (-0.462,-0.229)
BERTscore -0.318 0.011 -0.722
(-0.359,-0.276) (-0.032,0.054) (-0.836,-0.609)
BARTscore-F1 0.744 0.457 -0.042
(0.705,0.784) (0.427,0.487) (-0.163,0.079)
BARTscore-precision 0.877 0.536 -0.023
(0.831,0.924) (0.501,0.571) (-0.159,0.113)
BARTscore-recall -0.015 -0.014 -0.046
(-0.018,-0.012) (-0.017,-0.011) (-0.051,-0.041)
LMExaminer -1.245 -0.466 -0.204
(-1.328,-1.162) (-0.550,-0.382) (-0.340,-0.068)
GEM -0.386 -0.238 -0.516
(-0.459,-0.314) (-0.318,-0.159) (-0.624,-0.408)
GEM-S -0.361 -0.195 -0.283
(-0.436,-0.287) (-0.277,-0.114) (-0.385,-0.182)
Table 12: Standardized Mean Differences (SMD) $d:=(\mu^{\prime}-\mu)/{\sigma_{\text{pooled}}}$ of scores after degradations of various evaluation metrics with 95% confidence intervals in parentheses. Significant score decreases ($p<0.05$) after degradations are highlighted in bold green, implying the evaluation metric can effectively penalize the degradation.
GEM demonstrates robustness to manipulation.
Similarly, we reuse the GPT-4o Rephrase and Llama-3.1 Rephrase manipulations to test whether our metric is robust against specific language styles. However, since Yelp restaurant reviews lack a structured format like ICLR peer reviews (as illustrated in Figure 4), it is unclear how to construct a suitable Meaningless Elongation manipulation.
Table 13 shows the standardized mean differences (SMD) with 95% confidence intervals for GEM and baseline metrics before and after these manipulations. The results remain similar to those observed in Section 4.3. GEM metrics remain robust against both types of manipulation, while LMExaminer shows a significant increase in scores after LLM rephrasing.
Evaluation Metric GPT-4o Rephrase Llama3.1 Rephrase
BLEU 0.100 0.046
(0.014,0.185) (-0.063,0.156)
ROUGE-L 0.020 0.083
(-0.037,0.076) (0.008,0.157)
BERTscore -0.016 -0.098
(-0.041,0.008) (-0.133,-0.064)
BARTscore-F1 0.316 0.425
(0.287,0.346) (0.371,0.478)
BARTscore-precision 0.371 0.508
(0.337,0.405) (0.445,0.571)
BARTscore-recall -0.012 -0.016
(-0.013,-0.010) (-0.018,-0.014)
LMExaminer 0.300 0.453
(0.240,0.361) (0.387,0.518)
GEM 0.021 0.010
(-0.035,0.076) (-0.055,0.075)
GEM-S 0.020 -0.012
(-0.037,0.077) (-0.080,0.055)
Table 13: Standardized Mean Differences (SMD) of scores $d:=(\mu^{\prime}-\mu)/{\sigma_{\text{pooled}}}$ after manipulations of various evaluation metrics with 95% confidence intervals in parentheses. Significant score increases ($p<0.05$) after manipulations are highlighted in bold red, implying the evaluation metrics are not robust against the manipulation.
A.5 Exploration of Properties of LLM Estimated Distribution
An interesting question raised by our anonymous reviewer is:
As we know $\Pr[x,y]=\Pr[x]\cdot\Pr[y|x]=\Pr[y]\cdot\Pr[x|y]$, does this symmetry also applies to the LLM estimated probability, i.e., $\Pr_{\operatorname{LLM}}[x]\cdot\Pr_{\operatorname{LLM}}[y|x]=\Pr_{\operatorname{LLM}}[y]\cdot\Pr_{\operatorname{LLM}}[x|y]$?
To answer this question, we randomly select 500 papers from ICLR 2023 and 1874 corresponding reviews in total. For each pair of reviews $x$, $y$ for the same paper, we use our method to compute $\log\Pr_{\operatorname{LLM}}[x|y]$, $\log\Pr_{\operatorname{LLM}}[y|x]$, $\log\Pr_{\operatorname{LLM}}[x]$, and $\log\Pr_{\operatorname{LLM}}[y]$. For those 5328 pairs, we compute the Spearman correlation coefficient between $\Pr_{\operatorname{LLM}}[x]\cdot\Pr_{\operatorname{LLM}}[x|y]:=\exp\left(\log\Pr_{\operatorname{LLM}}[x|y]+\log\Pr_{\operatorname{LLM}}[y]\right)$ and $\Pr_{\operatorname{LLM}}[y]\cdot\Pr_{\operatorname{LLM}}[y|x]:=\exp\left(\log\Pr_{\operatorname{LLM}}[y|x]+\log\Pr_{\operatorname{LLM}}[x]\right)$. The Spearman correlation coefficient is 0.943, which indicates that they are highly positively correlated, despite minor differences. Potential noise in the estimation process may account for these differences.
The results revealed a high positive correlation between these two values, although they were not precisely equal. Despite this difference, the empirical results in our paper indicate that this does not negatively impact the desired properties, including the accuracy and manipulation resistance.